OpenStudy (lovechrist):

The information given represents lengths of sides of a right triangle with c the hypotenuse. Find the correct missing length to the nearest hundredth. A calculator may be helpful. a = ?, b = 11, c = 17 A. 12.75 B. 13.20 C. 12.96 D. 13.12

OpenStudy (aveline):

Wrong section o3o a^2+b^2=c^2 Apply this equation

OpenStudy (lovechrist):

?

OpenStudy (aveline):

What's the square root of 17^2-11^2?

OpenStudy (lovechrist):

-104?

OpenStudy (lovechrist):

-104?

OpenStudy (aveline):

289-121=?